# TextCrafter: Accurately Rendering Multiple Texts in Complex Visual Scenes

### v2

@@ Abstract
Figure 1: TextCrafter enables precise multi-region visual text rendering, addressing the challenges of long, small-size,various numbers, symbols and styles in visual text generation. We illustrate the comparisons among TextCrafter with three state-of-the-art models, i.e., FLUX, TextDiffuser-2 and 3DIS.
@@ 1 Introduction
In recent years, diffusion models Ho et al. (2020); Podell et al. (2023); Rombach et al. (2022); Ramesh et al. (2021, 2022); Zhao et al. (2024); Dhariwal and Nichol (2021); Chen et al. (2023) have emerged as the forefront technology in visual image generation. Recent studies, represented by general-purpose text-to-image models such as Flux BlackForest (2024) and SD 3.5 Esser et al. (2024), have acquired the ability to render simple visual text through large-scale pretraining. However, when confronted with complex real-world visual text scenarios, these models often struggle with critical challenges such as text distortion, omission, and blurriness. These shortcomings significantly hinder their practical applicability in real-world use cases. Achieving precise, legible, and contextually accurate visual text rendering in complex scenes remains essential yet highly challenging.
@@
Recent studies in multi-instance control Feng et al. (2022); Zhou et al. (2024); Chen et al. (2024b); Zhou et al. (2025) primarily focus on generating multiple instances within a single image simultaneously, striving to achieve precise regulation over factors such as quantity, position and attributes. However, a crucial yet often overlooked category in the visual domain is visual text. Text is an indispensable component of the real world, distinguished by its unparalleled complexity and delicacy compared to conventional objects, flora, or fauna. Unlike these entities, text possesses an exceptionally intricate internal structure, where even minor alterations can cause significant changes in visual appearance, potentially leading to misrecognition or complete illegibility. Existing multi-instance approaches, such as MIGC Zhou et al. (2024) and 3DISZhou et al. (2025), primarily focus on sentence-level information and interaction, offering a limited understanding of text at a finer granularity. Moreover, RPG Yang et al. (2024a) and RAG-Diffusion Chen et al. (2024b) adjust instance structures during attention fusion but fall short of the precision and stability needed for visual text rendering.
@@
Recent studies on visual text rendering primarily focus on accurately generating a single text region. For instance, methods like AnyText Tuo et al. (2024), Glyph-byT5 Liu et al. (2025), SceneTextGen Zhangli et al. (2024), Diff-Text Zhang et al. (2024b), and TextDiffuser Chen et al. (2024a) often employ specialized modules to encode visual text representations and condition the diffusion model accordingly. However, these methods face two major limitations: 1) They rely on predefined rules to synthesize visual text training datasets. Due to the need for highly accurate text annotations, manual verification is often required, making the creation of complex visual text datasets both labor-intensive and costly. 2) Most approaches depend on fine-tuned text encoders (e.g., Glyph-by-T5) or trained conditional control encoders (e.g., AnyText, Diff-Text, TextDiffuser, SceneTextGen) to facilitate text generation. In complex multi-text generation, however, these approaches often cause interference among the control features of different targets, making it difficult to balance fine-grained local control with global consistency across multiple texts.
@@
This paper explores the task of Complex Visual Text Generation (CVTG), which aims to generate intricate text content distributed across different regions. To tackle the challenges of CVTG, we propose TextCrafter, a novel training-free framework for rendering multiple texts in complex visual scenes. TextCrafter enhances multi-text control, ensuring accurate text presentation and logical layout. Our framework comprises three key stages: 1) Instance Fusion: This stage strengthens the connection between visual text and its corresponding carrier, ensuring strong consistency between the text content and its surrounding environment. It effectively prevents text from appearing in incorrect positions. 2) Region Insulation: Leveraging the positional priors of the pre-trained DiT model Peebles and Xie (2023) (e.g., FLUX), this step initializes the layout information for each text instance while separating and denoising text prompts across different regions. This process prevents early interference between text areas, reducing confusion and the risk of content omission in multi-text scenarios. 3) Text Focus: An attention control mechanism is introduced to enhance the attention maps of visual text, refining the fidelity of text rendering. This stage ensures textual accuracy and addresses issues of blurriness, particularly in smaller text. We present some examples in Figure 1. Our contributions are summarized as follows:
@@
Figure 2: Overview of our TextCrafter. TextCrafter consists of three steps. (a) Instance Fusion: Strengthen the connection between visual text and its corresponding carrier. (b) Region Insulation: Leverage the positional priors of the pre-trained DiT model to initialize the layout information for each text instance while separating and denoising text prompts across different regions. (c) Text Focus: Enhance the attention maps of visual text, refing the fidelity of text rendering.
@@ 2 Related Works
Attend-and-Excite Chefer et al. (2023) and BOX-Diffusion Xie et al. (2023) guide pre-trained diffusion models to align generated instances with prompts. GLIGEN Li et al. (2023), MIGC Zhou et al. (2024), and CreativeLayout Zhang et al. (2024a) enhance control by incorporating bounding box conditions into trainable layers. RPG Yang et al. (2024a) and RAG-Diffusion Chen et al. (2024b) break the generation process into regional tasks for compositional generation. However, these methods often overlook a critical element: visual text. As a key component of visual scenes, accurate text significantly enhances the realism of generated images.
@@
DiffSTE Ji et al. (2023), Glyph-ByT5 Liu et al. (2025), and UDiffText Zhao and Lian (2025) use character-level encoders to incorporate word appearance into embeddings. TextDiffuser Chen et al. (2024a) and GlyphControl Yang et al. (2024b) generate text layout masks for injection into the latent space. Approaches like Diff-Text Zhang et al. (2024b) improve accuracy with attention map restrictions. GlyphDraw Ma et al. (2023) uses two models for text position prediction and rendering. TextDiffuser-2 Chen et al. (2025) leverages large language models for layout planning. However, these approaches often optimize text rendering for single region or on individual object, leading to poor generalization in complex visual text scenes.
@@
Earlier works like GlyphControl introduced SimpleBench and CreativeBench Yang et al. (2024b), but these fixed-template datasets offer limited diversity and contain only single-word scenes. MARIO-Eval Chen et al. (2024a) suffers from low-quality prompts and unclear semantics. DrawTextExt Creative Ma et al. (2023) collects prompts for generating natural scenes but still lack sufficient coverage. AnyText-benchmark Tuo et al. (2024) allows multi-word prompts but treats each word separately and appends them directly to the caption, limiting data distribution. Existing visual text benchmarks focus on text generation within a single object or location, falling short of capturing the complexity of real-world scenes. Thus, we propose a benchmark CVTG-2K to incorporate diverse positions, quantities, lengths, and attributes for comprehensively evaluate models in complex visual text scenarios.
@@ 3.1 Overview
In CVTG, the user provides a global prompt $P$ that contains multiple descriptions of visual texts $D=\{d_{1},d_{2},\dots,d_{n}\}$, where each description includes the main body of the visual text and other descriptors such as the position and attributes of the visual text. The main body of the visual text is defined as $VT=\{vt_{1},vt_{2},\dots,vt_{n}\}$, where each $vt_{i}$ corresponds to all of its descriptors $d_{i}$. The goal of the model is to generate an image based on the provided prompt $P$, where each visual text $vt_{i}$ is presented with its corresponding description $d_{i}$.
@@
When handing complex prompts, diffusion models encounter interference among visual texts. This interference leads to several degradation issues. $1$) Text confusion: Visual texts (e.g., visual text $vt_{i}$ and visual text $vt_{j}$) intertwine, leading to the generation of erroneous character or word. $2$) Text omission: Only visual texts related to description $d_{i}$ are generated, neglecting the visual text $vt_{j}$ in description $d_{j}$. $3$) Text blurriness: Visual text in description $d_{i}$ is small in scale and fails to draw sufficient attention, causing $vt_{i}$ to appear blurred in generated image.
@@
Current text-to-image models excel at handing individual visual text in visual generation. By breaking down CVTG into simpler sub-tasks, overall complexity is reduced. Solving these sub-tasks separately and combining their solutions effectively accomplishes the CVTG task. This approach directs the model to progressively shift focus from the entire task to individual sub-tasks and then to specific visual texts, enabling precise rendering. Based on this coarse-to-fine strategy, we present TextCrafter, as shown in Figure 2, with detailed explanations provided in the following sections.
@@ 3.2 Instance Fusion
It is crucial to prevent visual texts from floating in the image and ensure proper positioning. We propose an instance fusion strategy that integrates the visual text’s content with its spatial carrier. Rather than using the text’s own embedding, we leverage embedding of the preceding quotation mark to establish spatial relationships. As illustrated in Figure 3, the preceding quotation mark in the attention map corresponds to the visual text it governs, indicating that it encapsulates complete information about the text. According to token additivity Hu et al. (2024), we incorporate the quotation mark’s embedding into the carrier’s embedding via weighted fusion to ensure precise alignment, where $\lambda$ controls the proportion of the addition. . Given a prompt $P$, after being processed by the T5 Raffel et al. (2020) text encoder $\tau_{\xi}$, the condition is expressed as $C=\tau_{\xi}(P)$. Each row of $C$ represents the token embedding corresponding to each token processed by T5.
Figure 4: For a pre-trained DiT model, only a few denoising steps are required to approximate the layout of the image and the relative positions of the main subjects. After 8 denoising steps, the layout closely resembles that of a full 50-step process, with subsequent steps primarily refining image details.
@@
CreativeBench 400 1.00 7.29 × ×
MARIOEval 5414 2.92 15.47 × ×
DrawTextExt 220 3.75 17.01 × ×
AnyText-benchmark 1000 4.18 21.84 × ×
@@
Input: The encoded embedding of $P$ and $D$ is denoted as $c$ and $\{\hat{c}^{1},\hat{c}^{2},\dots,\hat{c}^{n}\}$
Parameter: the proportion of the addition $\lambda$, the $r$ steps of region insulation, the pre-generation steps $\tau$
Output: A generated latent $x_{0}$
1: {Step 1 Instance Fusion}
2:$c=\text{InstanceFusion}(c,\lambda)$
3: {Step 2 Region Insulation}
4:$z_{t}\sim\mathcal{N}(0,1)$
5:for $t=\tau,\tau-1,\dots,1$ do
6:    $z_{t-1},A_{t}\leftarrow\text{DiT}(z_{t},c,t)$
7:end for
8:for $i=1,2,\dots,n$ do
9:    $p_{i}\leftarrow(0,0)$
10:    $p_{\text{max},i}\leftarrow\arg\max_{p_{i}}A_{t}(p_{i},i)$
11:end for
12:$bbx_{1},bbx_{2},\dots,bbx_{n}\leftarrow\text{MILP}(p_{\text{max},1},p_{\text{max},2},\dots,p_{\text{max},n})$
13:$x_{T}\sim\mathcal{N}(0,1)$
14:for $i=1,2,\dots,n$ do
15:    $\hat{x}_{t}^{i}\leftarrow\text{Initialize}(\mathcal{N}(0,1),bbx_{i})$
16:end for
17:for $t=T,T-1,\dots,T-r+1$ do
18:    $x_{t-1}\leftarrow x_{t}-\epsilon_{\theta}(x_{t},c)$
19:    for $i=1,2,\dots,n$ do
20:       $\hat{x}_{t-1}^{i}\leftarrow\hat{x}_{t}^{i}-\epsilon_{\theta}(\hat{x}_{t}^{i},\hat{c}^{i})$
21:       $x_{t-1}\leftarrow\text{Reinsert}(x_{t-1},\hat{x}_{t-1}^{i},bbx_{i})$
22:    end for
23:end for
24: {Step 3 Text Focus}
25:for $t=T-r,T-r-1,\dots,1$ do
26:    for $AttentionLayer_{i}\in\epsilon_{\theta}$ do
27:       $M_{t}=Q_{t}^{i}\times K_{t}^{i}$
28:       $M_{t}\leftarrow\text{Focus}(M_{t})$
29:       $Attention_{t}^{i}\leftarrow\text{softmax}(M_{t}/\sqrt{d})\times V_{i}$
30:    end for
31:    $x_{t-1}\leftarrow x_{t}-\epsilon_{\theta}(x_{t},c)$
32:end for
33:return $x_{0}$
Word Accuracy↑
Model 2 regions 3 regions 4 regions 5 regions average NED↑ CLIPScore↑
SD3.5 Large (ICML 2024) 0.7293 0.6825 0.6574 0.5940 0.6548 0.8470 0.7797
FLUX.1 dev (2024) 0.6089 0.5531 0.4661 0.4316 0.4965 0.6879 0.7401
AnyText (ICLR 2024) 0.0513 0.1739 0.1948 0.2249 0.1804 0.4675 0.7432
TextDiffuser-2 (ECCV 2024) 0.5322 0.3255 0.1787 0.0809 0.2326 0.4353 0.6765
RAG-Diffusion (Arxiv 2024) 0.4388 0.3316 0.2116 0.1910 0.2648 0.4498 0.7797
3DIS (ICLR 2025) 0.4495 0.3959 0.3880 0.3303 0.3813 0.6505 0.7767
TextCrafter (Ours) 0.7628 0.7628 0.7406 0.6977 0.7370 0.8679 0.7868
Table 2: Quantitative results on the proposed CVTG-2K. The results highlight TextCrafter’s superiority across three metrics.
@@ 3.3 Region Insulation
In CVTG, multiple texts often influence each other’s glyphs and layout Zheng et al. (2023). To simplify, we decouple them into independent instances. Specifically, we assign a rectangular bounding box to each visual text to represent its layout, initialized based on the positional preferences of the pre-trained DiT model. This allows us to effectively isolate the texts, preventing interference during generation. Instead of manually specifying layouts or relying on large language models, we leverage the model’s learned positional knowledge Tumanyan et al. (2023), as enforcing a fixed layout may compromise quality. As shown in Figure 4, during the early denoising steps, we capture attention maps and identify the points of maximum attention for each visual text. This pre-generation process guides layout initialization. Assume that the attention map obtained during this process is $A_{t}$, where $t$ represents the current step, and $p$ stands for pixel:
@@
$x_{t-1}$ $=x_{t}-\epsilon_{\theta}(x_{t},c)$
$\hat{x}_{t-1}^{i}=\hat{x}_{t}^{i}-\epsilon_{\theta}(\hat{x}_{t}^{i},\hat{c}^{i}),$ (3)

### v3

@@ Abstract
Figure 1: TextCrafter enables precise multi-region visual text rendering, addressing the challenges of long, small-size, various numbers, symbols and styles in visual text generation. We illustrate the comparisons among TextCrafter with three state-of-the-art models, i.e., FLUX BlackForest [1], TextDiffuser-2 Chen et al. [5] and 3DIS Zhou et al. [45].
@@ 1 Introduction
In recent years, diffusion models  [13, 27, 30, 28, 29, 40, 8, 6, 17] have emerged as the forefront technology in visual image generation. Recent studies, represented by general-purpose text-to-image models such as Flux [1] and SD 3.5 [9], have acquired the ability to render simple visual text through large-scale pretraining. However, when confronted with complex real-world visual text scenarios, these models often struggle with critical challenges such as text distortion, omission, and blurriness. These shortcomings significantly hinder their practical applicability in real-world use cases. Achieving precise, legible, and contextually accurate visual text rendering in complex scenes remains essential yet highly challenging.
@@
Recent studies in multi-instance control [10, 43, 7, 45, 18] primarily focus on generating multiple instances within a single image simultaneously, striving to achieve precise regulation over factors such as quantity, position and attributes. However, a crucial yet often overlooked category in the visual domain is visual text. Text is an indispensable component of the real world, distinguished by its unparalleled complexity and delicacy compared to conventional objects, flora, or fauna. Unlike these entities, text possesses an exceptionally intricate internal structure, where even minor alterations can cause significant changes in visual appearance, potentially leading to misrecognition or complete illegibility. Existing multi-instance approaches, such as MIGC [43] and 3DIS [45], primarily focus on sentence-level information and interaction, offering a limited understanding of text at a finer granularity. Moreover, RPG [35] and RAG-Diffusion [7] adjust instance structures during attention fusion but fall short of the precision and stability needed for visual text rendering.
@@
Recent studies on visual text rendering primarily focus on accurately generating a single text region. For instance, methods like AnyText [32], Glyph-byT5 [23], SceneTextGen [39], Diff-Text [38], and TextDiffuser [4] often employ specialized modules to encode visual text representations and condition the diffusion model accordingly. However, these methods face two major limitations: 1) They rely on predefined rules to synthesize visual text training datasets. Due to the need for highly accurate text annotations, manual verification is often required, making the creation of complex visual text datasets both labor-intensive and costly. 2) Most approaches depend on fine-tuned text encoders (e.g., Glyph-by-T5) or trained conditional control encoders (e.g., AnyText, Diff-Text, TextDiffuser, SceneTextGen) to facilitate text generation. In complex multi-text generation, however, these approaches often cause interference among the control features of different targets, making it difficult to balance fine-grained local control with global consistency across multiple texts.
@@
This paper explores the task of Complex Visual Text Generation (CVTG), which aims to generate intricate text content distributed across different regions. To tackle the challenges of CVTG, we propose TextCrafter, a novel training-free framework for rendering multiple texts in complex visual scenes. TextCrafter enhances multi-text control, ensuring accurate text presentation and logical layout. Our framework comprises three key stages: 1) Instance Fusion: This stage strengthens the connection between visual text and its corresponding carrier, ensuring strong consistency between the text content and its surrounding environment. It effectively prevents text from appearing in incorrect positions. 2) Region Insulation: Leveraging the positional priors of the pre-trained DiT model [26] (e.g., FLUX), this step initializes the layout information for each text instance while separating and denoising text prompts across different regions. This process prevents early interference between text areas, reducing confusion and the risk of content omission in multi-text scenarios. 3) Text Focus: An attention control mechanism is introduced to enhance the attention maps of visual text, refining the fidelity of text rendering. This stage ensures textual accuracy and addresses issues of blurriness, particularly in smaller text. We present some examples in Figure 1. Our contributions are summarized as follows:
@@
Figure 2: Overview of our TextCrafter. TextCrafter consists of three steps. (a) Instance Fusion: strengthen the connection between visual text and its corresponding carrier. (b) Region Insulation: leverage the positional priors of the pre-trained DiT model to initialize the layout information for each text instance while separating and denoising text prompts across different regions. (c) Text Focus: enhance the attention maps of visual text, refining the fidelity of text rendering.
@@ 2 Related Works
Attend-and-Excite [3] and BOX-Diffusion [34] guide diffusion models to align with prompts. GLIGEN [21], MIGC [43], and CreativeLayout [37] incorporate bounding boxes into trainable layers. RPG [35], RAG-Diffusion [7] and DreamRenderer [44] decompose generation into regional tasks. However, these methods overlook visual text, a critical element for image realism.
@@
DiffSTE [19], Glyph-ByT5 [23], and UDiffText [41] use character-level encoders for word appearance. TextDiffuser [4] and GlyphControl [36] generate text layout masks. Diff-Text [38] restricts attention maps. GlyphDraw [24] utilizes dual models for positioning and rendering. TextDiffuser-2 [5] leverages LLMs for layout planning. These approaches optimize for single region or individual object, limiting generalization in complex scenes.
@@
GlyphControl introduced SimpleBench and CreativeBench [36], but these contain only single-word scenes. MARIO-Eval [4] has low-quality prompts. DrawTextExt Creative [24] lacks sufficient coverage. AnyText-benchmark [32] treats words separately, limiting distribution. Existing benchmarks focus on a single object or location, failing to capture real-world complexity. Thus, we propose CVTG-2K to evaluate models with diverse positions, quantities, lengths, and attributes in complex visual text scenarios.
@@
Figure 4: Motivation for using a layout optimizer. LLM/MLLM-based layout designers frequently generate overlapping bounding boxes, which do not align with the layout preferences of Diffusion models. Our proposed layout optimizer addresses this limitation without requiring additional fine-tuning.
@@ 3.1 Overview
In CVTG, users provide a global prompt $P$ containing multiple visual text descriptions $D=\{d_{1},d_{2},\dots,d_{n}\}$, where each description includes the text’s main body and descriptors (position, attributes). The main text bodies are defined as $VT=\{vt_{1},vt_{2},\dots,vt_{n}\}$, with each $vt_{i}$ corresponding to descriptors $d_{i}$. The model must generate an image from prompt $P$ where each $vt_{i}$ appears with its corresponding description $d_{i}$.
@@
Complex prompts cause interference among visual texts, leading to three degradation issues: $1$) Text confusion: Visual texts $vt_{i}$ and $vt_{j}$ intertwine, generating erroneous characters/words. $2$) Text omission: Only text from description $d_{i}$ appears, neglecting $vt_{j}$ from $d_{j}$. $3$) Text blurriness: Small-scale text in $d_{i}$ receives insufficient attention, appearing blurred in the generated image.
@@
Current text-to-image models excel with individual visual texts. By decomposing CVTG into simpler sub-tasks, overall complexity decreases. Solving these sub-tasks separately and combining their solutions accomplishes the CVTG task. This approach shifts focus progressively from the entire task to sub-tasks then to specific visual texts, enabling precise rendering. Based on this coarse-to-fine strategy, we present TextCrafter (Figure 2), detailed in subsequent sections.
@@ 3.2 Instance Fusion
It is crucial to prevent visual texts from floating in the image and ensure proper positioning. We propose an instance fusion strategy that integrates the visual text’s content with its spatial carrier. Rather than using the text’s own embedding, we leverage embedding of the preceding quotation mark to establish spatial relationships. As illustrated in Figure 3, the preceding quotation mark in the attention map corresponds to the visual text it governs, indicating that it encapsulates complete information about the text. According to token additivity [14], we incorporate the quotation mark’s embedding into the carrier’s embedding via weighted fusion to ensure precise alignment, where $\lambda$ controls the proportion of the addition.
Figure 5: For a pre-trained DiT model, only a few denoising steps are required to approximate the image layout and subject positions. After 8 steps, the layout closely resembles that of a full 50-step process, with subsequent iterations primarily refining details.
@@
CreativeBench 400 1.00 7.29 $\times$ $\times$
MARIOEval 5414 2.92 15.47 $\times$ $\times$
DrawTextExt 220 3.75 17.01 $\times$ $\times$
AnyText-benchmark 1000 4.18 21.84 $\times$ $\times$
@@
1:  Input: The encoded embedding of a global prompt $P$ is denoted as $c$. The encoded embedding of multiple descriptions $D$ is denoted as $\{\hat{c}^{1},\hat{c}^{2},\dots,\hat{c}^{n}\}$. 2:  Parameter: the proportion of the addition $\lambda$, the $r$ steps of Region Insulation, the pre-generation steps $\tau$. 3:  Output: A generated latent $x_{0}$ 4:  {Step 1: Instance Fusion} 5:  $c=\text{InstanceFusion}(c,\lambda)$ 6:  {Step 2: Region Insulation} 7:  $z_{t}\sim\mathcal{N}(0,1)$ 8:  for $t=\tau,\tau-1,\dots,1$ do 9:     $z_{t-1},A_{t}\leftarrow\text{DiT}(z_{t},c,t)$ 10:  end for 11:  for $i=1,2,\dots,n$ do 12:     $p_{i}\leftarrow(0,0)$ 13:     $p_{\text{max},i}\leftarrow\arg\max_{p_{i}}A_{t}(p_{i},i)$ 14:  end for 15:  $bbx_{1},\dots,bbx_{n}\leftarrow\text{MILP}(p_{\text{max},1},\dots,p_{\text{max},n})$ 16:  $x_{T}\sim\mathcal{N}(0,1)$ 17:  for $i=1,2,\dots,n$ do 18:     $\hat{x}_{t}^{i}\leftarrow\text{Initialize}(\mathcal{N}(0,1),bbx_{i})$ 19:  end for 20:  for $t=T,T-1,\dots,T-r+1$ do 21:     $x_{t-1}\leftarrow x_{t}-\epsilon_{\theta}(x_{t},c)$ 22:     for $i=1,2,\dots,n$ do 23:        $\hat{x}_{t-1}^{i}\leftarrow\hat{x}_{t}^{i}-\epsilon_{\theta}(\hat{x}_{t}^{i},\hat{c}^{i})$ 24:        $x_{t-1}\leftarrow\text{Reinsert}(x_{t-1},\hat{x}_{t-1}^{i},bbx_{i})$ 25:     end for 26:  end for 27:  {Step 3: Text Focus} 28:  for $t=T-r,T-r-1,\dots,1$ do 29:     for $\text{AttentionLayer}_{i}\in\epsilon_{\theta}$ do 30:        $M_{t}=Q_{t}^{i}\times K_{t}^{i}$ 31:        $M_{t}\leftarrow\text{Focus}(M_{t})$ 32:        $\text{Attention}_{t}^{i}\leftarrow\text{softmax}(M_{t}/\sqrt{d})\times V_{i}$ 33:     end for 34:     $x_{t-1}\leftarrow x_{t}-\epsilon_{\theta}(x_{t},c)$ 35:  end for 36:  return  $x_{0}$
Table 2: Quantitative results on the proposed CVTG-2K. The results highlight TextCrafter’s superiority across five metrics.
Model Word Accuracy↑ NED↑ CLIPScore↑ VQAScore↑ Aesthetics↑
FLUX.1 dev (2024) 0.4965 0.6879 0.7401 0.8886 5.91
SD3.5 Large (ICML 2024) 0.6548 0.8470 0.7797 0.9297 5.56
AnyText (ICLR 2024) 0.1804 0.4675 0.7432 0.6935 4.53
TextDiffuser-2 (ECCV 2024) 0.2326 0.4353 0.6765 0.5627 4.51
3DIS (ICLR 2025) 0.3813 0.6505 0.7767 0.8684 4.86
TextCrafter (FLUX) 0.7370 0.8679 0.7868 0.9140 5.70
TextCrafter (SD3.5) 0.7600 0.9038 0.8023 0.9299 5.45
@@ 3.3 Region Insulation
In CVTG, multiple texts often influence each other’s glyphs and layout [42]. To simplify, we decouple them into independent instances. Specifically, we assign a rectangular bounding box to each visual text to represent its layout, initialized based on the positional preferences of the pre-trained DiT model. This allows us to effectively isolate the texts, preventing interference during generation. Instead of manually specifying layouts or relying on LLMs, we leverage the diffusion model’s learned positional knowledge [31, 16]. Figure 4 demonstrates a notable limitation of LLMs/MLLMs: despite dedicated fine-tuning, the layouts they provide often exhibit overlapping bounding boxes and do not conform to the layout preferences inherent in Diffusion models. As shown in Figure 5, during the early denoising steps, we capture attention maps and identify the points of maximum attention for each visual text. This pre-generation process guides layout initialization. Assume that the attention map obtained during this process is $A_{t}$, where $t$ represents the current step, and $p$ stands for pixel:
@@
$x_{t-1}=x_{t}-\epsilon_{\theta}(x_{t},c),\quad\hat{x}_{t-1}^{i}=\hat{x}_{t}^{i}-\epsilon_{\theta}(\hat{x}_{t}^{i},\hat{c}^{i})$ (3)